40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  TCS elog  Not logged in ELOG logo
Message ID: 234     Entry time: Wed Jul 24 16:25:13 2019
Author: Edita Bytyqi 
Type: Lab Infrastructure 
Category: Optics 
Subject: Updated 2-lens setup 

The previous 2-lens setup focused the beam to a tight spot, however due to the divergence angle of the laser beam, a significant amount of power was not being captured by the fiirst lens at a distance of 40 cm from the source. The divergence angle seems to be bigger than 0.06 by a factor of 2, so a f = 20 cm lens was used to collimate the beam and a f = 30 cm lens was used to focus it. A mirror was used to reflect the beam, so we obtain steering control. Additionally, the focusing lens was placed on a small 1-axis stage in order to control the distance of the lens from the CCD, providing control over the focused beam size.

Note: The 30 cm lens was cleaned with methanol, however it still has some residue on the surface. The beam imaged to the Harrtmann Sensor looks good, however the lens will be cleaned by using a different solvent or replaced by a different 30 cm lens. The 3 lenses at the edge of the box will stay inside in order to prevent contamination, however they will not be used in the design.

Attachment 1: 20190724_161926.jpg  4.442 MB  Uploaded Wed Jul 24 17:34:14 2019  | Show | Show all
ELOG V3.1.3-